OpenStudy (anonymous):

To solve by completing the square, what value should you add to each side of the equation. x^2+16x=-4

OpenStudy (anonymous):

\[(\frac{ b }{ 2 })^{2}\] use this. b in this equation is 16. that will give you the value.

OpenStudy (anonymous):

Completing the square hinges on the algebraic property of raising a binomial to the power of \(2\). Observe:$$(a+b)^2=(a+b)(a+b)=a(a+b)+b(a+b)=a^2+ab+ab+b^2=a^2+2ab+b^2$$ In our case, we observe we have \(x^2+16x\) corresponding to \(a^2+2ab\), so \(x\) is our \(a\) and therefore \(b\) must be \(\frac12(16)=8\). Notice all we're missing, then, is our final term \(b^2\). Since we deduced \(b=8\), we add \(b^2=64\) to our equation to yield:$$x^2+16x+64=-4+64=60\\(x+8)^2=6$$

OpenStudy (anonymous):

So 8^2 = 64?

OpenStudy (anonymous):

\(60\)*

OpenStudy (anonymous):

so I would need to add 60 to each side of the equation?

OpenStudy (anonymous):

Yes. that's it

OpenStudy (anonymous):

16/2 = 8 8*8 = 64

OpenStudy (anonymous):

@PECKE no you add \(64\), I was merely correcting my last line to \((x+8)^2=60\).

OpenStudy (anonymous):

ahh okay, phew think I got it all Thanks